OpenStudy (anonymous):

Rewrite using standard notation: 3.58 x 10-2

OpenStudy (mathstudent55):

10^(-2) = 1/(10^2) = 1/100 = 0.01 Just multiply 3.58 by 0.01, or divide 3.58 by 100. Both will give you the same answer.